STATISTICS. Suppose two numbers, x and y, generated at random, where 0 < x < 5 and 0 < y <10. What is the probability that the sum is less than or equal to 2? [hint: x and y are not discrete variables. Use the concept of geometric probability]
Answer by greenestamps(13200)



The "sample space" is the region in the xy plane bounded by x=0 and x=5 and by y=0 and y=10; the area of that region is 5*10=50.

The outcomes we want are those bounded by x=0, y=0, and x+y=2. The area of that triangular region is (1/2)(2)(2) = 2.

The probability that the sum of the two numbers is less than or equal to 2 is 2/50 = 1/25.